OpenStudy (anonymous):

Laura has 23 coins in a jar. The jar contains only dimes and quarters. If the coins are worth $4.85, how many dimes does Laura have?

OpenStudy (anonymous):

let x be the number of dimes, and y be the number of quarters. Since there are 23 coins, we know that: \[x+y = 23\]Then, because the value of those coins is 4.85, we know that; \[10x+25y = 485\] Now we just have to solve this system of equations.

OpenStudy (anonymous):

Let x be the number of dimes and y be the number of quarters So, x+y = 23 0.1x+0.25y = 4.85 10x+25y = 485 Solce simultaneously, to get x=6 and y = 17

OpenStudy (anonymous):

You could multiply the first equation by 10 to get: \[10x+10y =230 \]Then subtract the two equations to get: \[15y = 255 \iff y = 17\] Now we know there are 17 quarters. So there must be 6 dimes (since there were 23 coins in the jar)

OpenStudy (amistre64):

23 coins dimes and quarters worth $4.85 d = (4.85 - 23(.25))/(.10-.25) = 6